OpenStudy (nathanjhw):

Let f be the function shown below, with domain the closed interval [0, 6]. The graph of f shown below is composed of two semicircles.

OpenStudy (nathanjhw):

h(x) = integral (0,2x-1) f (t) dt

OpenStudy (nathanjhw):

A. Determine the domain of h(x). B. Find h ' (5/2) C. At what x is h(x) a maximum?

OpenStudy (nathanjhw):

@phi

OpenStudy (phi):

can you post a png or pdf version rather than a docx?

OpenStudy (nathanjhw):

Yeah one sec.

OpenStudy (phi):

I'm thinking you find the domain of h(x) using 2x-1 = 0 and 2x-1 = 6 to get [0.5, 3.5]

OpenStudy (nathanjhw):

That seems right to me. I already did A and got the same answer. Just wanted to double check.

OpenStudy (nathanjhw):

@phi How would I do B and C?

OpenStudy (phi):

B. Find h ' (5/2) they want the value of the derivative of h(x) at x=5/2 h(x) is the integral of f(2x-1) , so the derivative h'(x) is f(2x-1)

OpenStudy (nathanjhw):

Is f(2x-1) the answer or do we need to do f(2x-1) = 5/2 or something else?

OpenStudy (phi):

you have f(t) plotted t= 2x-1 plug in x=5/2 to find what t is. then look for f(t) on the graph

OpenStudy (nathanjhw):

t= 2(5/2) -1 t = 5 -1 t=4

OpenStudy (phi):

now look at f(4) what is its value?

OpenStudy (nathanjhw):

0

OpenStudy (phi):

that is the answer to B.

OpenStudy (phi):

C. At what x is h(x) a maximum? do you remember how to determine the max of a curve?

OpenStudy (nathanjhw):

Before we move on, I checked my answer for B really quickly on wolfram. https://www.wolframalpha.com/input/?i=h%28x%29%3D+integral+%280%2C+2x-1%29+f%28t%29+dt+at+x+%3D5%2F2

OpenStudy (nathanjhw):

Do I write is as the integral or f(4)=0

OpenStudy (phi):

B. Find h ' (5/2) h(x) is the integral of f(t) h'(x) is f(t) in other words, you want the value of f(4) in this case.

OpenStudy (nathanjhw):

Alright, I understand now.

OpenStudy (nathanjhw):

So for C is ax2 + bx + c = 0 the formula?

OpenStudy (phi):

I think there is one subtlety and h'(x) = f(2x-1) * d/dx(2x-1) or \[ \frac{dh}{dx} = 2f(t)\] but when f(t) is 0, we still get 0

OpenStudy (phi):

For C C. At what x is h(x) a maximum? this is calculus. what do you recall when you hear min or max?

OpenStudy (nathanjhw):

Min is the lowest point and maximum is the highest point.

OpenStudy (phi):

in terms of derivatives?

OpenStudy (nathanjhw):

I believe it's: less than 0, it is a local maximum greater than 0, it is a local minimum equal to 0, then the test fails

OpenStudy (phi):

If you are at the peak of a smooth curve, what is the slope of the tangent line at that point?

OpenStudy (nathanjhw):

0

OpenStudy (phi):

and how do you find the slope of the curve at some point ?

OpenStudy (nathanjhw):

(y2-y1)/(x2-x1)

OpenStudy (phi):

in calculus you do that for two points that are *very* close together assuming y1= f(x) and y2= f(x+h): \[ \lim_{h\rightarrow 0} \frac{f(x+h) - f(h)}{(x+h)- x }\]

OpenStudy (phi):

Have you seen that before?

OpenStudy (phi):

\[ f'(x) = \lim_{h\rightarrow 0} \frac{f(x+h) - f(h)}{h } \]

OpenStudy (phi):

The idea is to find the derivative of h(x): find h'(x) and set it equal to 0 to find the min or max points. in your problem h'(x) = 2 f(t) 2 f(t) = 0 f(t)=0 find the t's where f(t) is 0 that happens at t=0, t=4 and t=6

OpenStudy (phi):

to decide if a point is a min or max, look at the second derivative. a positive second derivative is a min we want a max, which means we want a negative second derivative.

OpenStudy (phi):

to find the second derivative look at the tangent lines to f(t) near t=0 The line slopes up and to the right in that region. the positive slope means min at t=4 the tangent line goes down and to the right. so that is our max at t=6, it looks like another min

OpenStudy (nathanjhw):

Thanks! You really helped me out!